# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v2

@@ III.2 Quantum chemistry matrices
We start with preliminary qualitative remarks on the suitability of HHL/Psi-HHL for carrying out LCC calculations, and to that end, discuss the growth of the $A$ matrix dimension and $\mathcal{\kappa}$ with system size. This in turn requires us to define system size. Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we take the approach of: a. choosing a few simple diatomic closed-shell molecules in their ground state and in their equilibrium bond lengths, b. picking a suitable single particle basis, c. deciding on the number of occupied orbitals, and d. systematically increasing the number of unoccupied orbitals (virtuals) for system size. We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24]. The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule. We pick the Sapporo double zeta basis sets for KH, RbH, and CsH (K, Rb, and Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [25], while for Li2, we choose the STO-6G basis [26]. We employ the $C_{2v}$ point group symmetry for KH, RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry. We do not include relativistic effects as our aim is to only compare HHL and Psi-HHL and not account for as many relevant physical effects as possible. We now proceed to discuss the dimension of the $A$ matrix and the condition number scaling in LCC for these chosen molecules.
@@ III.2.1 Dimension of the $A$ matrix
Since the LCC matrix is not guaranteed to be of dimension $2^{N}\equiv n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$. We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix. We note that $m$, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for an LCC calculation, which in turn is decided by choice of single particle basis and active space size. We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$. The $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential. $\mathcal{E}$ refers to the excitation rank. For example, $\mathcal{E}=2$ is the LCCSD approximation, where S stands for single excitations and D for double excitations. For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix. For the purposes of this work, we study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number, all within the LCCSD approximation. The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector for our computations are discussed in Section A2.A of the Appendix.
@@ III.2.2 Condition number scaling in LCC
Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered. We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$. We increase the size of the active space (via $n_{v}$), and thus $m$ in a systematic manner, by carrying out an orbital character analysis whose details we explain in Section A2.B of the Appendix and also accompanying Table A10. Furthermore, for this analysis, we consider KH, RbH and CsH, but not Li2, as the system is too small (in the STO-6G basis) to carry out a $\mathcal{\kappa}$ versus $m$ analysis. We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase the number of virtuals. Figure 6 presents our results. We see that the condition number varies as a polylogarithmic function of system size. This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A11 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for KH, RbH, and CsH. The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix. We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration. The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the Hartree-Fock (HF) configuration). Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value. Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size.
We also observe in each sub-figure (that is, for a fixed number of occupied orbitals), for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH. This trend can be explained qualitatively using chemical intuition, by looking at the considered orbitals for the (6, 6) case as a representative example. Each eigenvalue of $A$ can be thought of as an excited state energy, given that $A$ is sparse and diagonally dominant. The highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively. It is clear that $2p$ of K has much lower orbital energy than the $3d$ orbital of Rb. Thus, we can expect that the $2p$longrightarrow virtual excitation in KH has a higher excitation energy associated with it than the 3d $\longrightarrow$ virtual excitation in RbH. Similarly, we can expect that the $4d$longrightarrow virtual excited state of CsH is lower in energy than the $3d$longrightarrow virtual excited state of RbH. In summary, we can expect that the excitation energies decrease monotonically as we go from KH to CsH. We combine this observation with our data from Table A11 where the maximum eigenvalue grows much faster than the decrease in minimum eigenvalue from KH to RbH to CsH for a given choice of $n_{o}$ and $n_{v}$, to infer that $\mathcal{\kappa}$ decreases monotonically from KH to CsH.
Furthermore, in Figure 6(c), KH’s condition number reaches a staggering 4000 with all the virtuals considered, whereas in Figures 6(a) and (b), $\mathcal{\kappa}$ is substantially lower. This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals, and KH has only 10 occupied orbitals and thus the active space contains the $1s$ orbital of K. We can invoke the fact that the orbital energy gap is larger between the inner s orbitals, for example, and the density of orbitals becomes higher in the vicinity of highest occupied molecular orbital and the lowest unoccupied molecular orbital. Thus, if we include inner core orbitals in our active space, we expect that the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger. In contrast, for the 6, 8, and the 10 occupied orbitals’ cases, RbH and CsH have comparable condition numbers. This can be attributed to the fact that the lowest orbital is always the $3d$ of Rb in all the three cases. Since the orbital character of the lowest orbital is the same among the three active spaces, they yield very similar condition numbers.
@@
Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for two representative molecules, via PFDs of Li2 (sub-figure (a)) and KH (sub-figure (b)) molecules versus number of shots (denoted in the figure as ‘No. of shots’). The centre of each unfilled circle represents the average PFD over a range of 30 repetitions. Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40). The zero PFD value is marked with a horizontal line for ease of visualization.
Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs (sub-figure (b)) vs bond lengths of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots) and those computed with the classical LCC value. The associated value of $\mathcal{\kappa}$ is shown just above or below each data point.
@@
We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider. However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation.
@@ III.2.3 HHL and Psi-HHL results
We begin by conducting a study similar to the one shown in Figure 5 for the toy matrices. It is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon. We next move to Figure 7 (with the accompanying data given in Table A12 of the Appendix), where we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions. For this analysis, we consider Li2 and KH. The details are:
@@
We pick these two systems, as in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300. Repeating the calculations for RbH and CsH may not give any new insights, and hence we restrict the results only to Li2 and KH. For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots. The data clearly shows that for Li2, the Psi-HHL approach always has a smaller spread in PFDs when compared to HHL. For KH, we observe that at 1000 shots, HHL fails, while Psi-HHL does not, but comes with large error bars. As we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller, but Psi-HHL consistently has smaller error bars than HHL.
@@
We now move to our next set of results, namely the PEC for the KH molecule. Since it is computationally not feasible to pick the larger system sizes from Figure 6, we choose the 6 occupied and 6 virtual orbitals case. This choice corresponds to a 26-qubit calculation with an $A$ matrix size of $256\times 256$. We consider geometries in the neighbourhood of the equilibrium bond length (3 geometries in steps of 0.47 Bohr on either side of the equilibrium bond length). With such a choice, the correlation energies required to be captured ranges from about 0.3 mHa (milli-Hartree) for 5.65 Bohr geometry to 3 mHa for the 2.82 Bohr geometry (the two extreme points in the 7 geometries considered), thus allowing us to check the performance of HHL and Psi-HHL across a range of correlation energy values. Calculating energies across a range of bond lengths further allows us to check across a (limited) range of condition numbers (the smallest being 365.40 and the largest 393.21) the behaviour of HHL and Psi-HHL. Each calculation involves 50000 shots and our reported results are averaged over 30 repetitions. For our Psi-HHL computations, we pick $\alpha$ of 60 degrees. Figure 8(a) (and the accompanying Table A13 of the Appendix) shows the results for $\Delta E$, which is the difference between HHL (Psi-HHL) total energy and and the classical LCC value for the considered geometries. We find that Psi-HHL performs better than HHL in predicting correlation energies, especially on the extreme ends of our PEC when $\mathcal{\kappa}$ values are relatively larger than for those points around the equilibrium geometry. When we examine Figure 8(b), we see that Psi-HHL performs consistently better than HHL. Furthermore, the spread in results across 30 repetitions follows a trend globally: it increases from 2.82 through 5.65 Bohr calculations. We recall that the correlation energies themselves decrease monotonically and by as much as an order of magnitude as we move from 2.82 Bohr to 5.65 Bohr geometries. Thus, although the condition numbers are comparable (for example, 374.93 and 371.38 for the 3.29 and the 4.71 Bohr cases respectively), since the quantity to be captured (the correlation energy) itself is much smaller for the latter, we need to supply more shots to restrict the spread in PFD across repeats. However, since we keep the number of shots fixed throughout the PEC, we observe a larger spread in PFDs. We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using HHL, with 1000 shots and 20 repeats. We average our results over 20 repetitions. $a$ is varied from 0.26 to 0.99 (below 0.26, we get nan). We observe the same trend as in Figure 8(b) (see Figure A10 of Appendix).

### v3

@@ III.2 Quantum chemistry matrices
We start with preliminary qualitative remarks on the suitability of HHL/Psi-HHL for carrying out LCC calculations, and to that end, discuss the growth of the $A$ matrix dimension and $\mathcal{\kappa}$ with system size. This in turn requires us to define system size. Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we follow the following steps: a. choosing a few simple diatomic closed-shell molecules in their ground state and in their equilibrium bond lengths, b. picking a suitable single particle basis, c. deciding on the number of occupied orbitals, and d. systematically increasing the number of unoccupied orbitals (virtuals) for system size. We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24]. The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule. We pick the Sapporo double zeta basis sets for KH, RbH, and CsH (K, Rb, and Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [25], while for Li2, we choose the STO-6G basis [26]. We employ the $C_{2v}$ point group symmetry for KH, RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry. We do not include relativistic effects as our aim is to only compare HHL and Psi-HHL and not account for as many relevant physical effects as possible. We now proceed to discuss the dimension of the $A$ matrix and the condition number scaling in LCC for these chosen molecules.
@@ III.2.1 Dimension of the $A$ matrix
Since the LCC matrix is not guaranteed to be of dimension $2^{N}\times 2^{N}\equiv n\times n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$. We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix. We note that $m$, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for an LCC calculation, which in turn is decided by choice of single particle basis and active space size. We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$. The number of rows (or columns) of the $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential. $\mathcal{E}$ refers to the excitation rank. For example, $\mathcal{E}=2$ is the LCCSD approximation, where S stands for single excitations and D for double excitations. For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially (for example, see Figure A11 in the Appendix). For the purposes of this work, we fix $n_{o}$ and increase $n_{v}$ for a given molecule, all within the LCCSD approximation. The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector for our computations are discussed in Section A2.A of the Appendix.
@@ III.2.2 Condition number scaling in LCC
Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the molecules that we consider. We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$. We increase the size of the active space (via $n_{v}$), and thus $m$ in a systematic manner, by carrying out an orbital character analysis whose details we explain in Section A2.B of the Appendix and also accompanying Table A13. Furthermore, for this analysis, we consider KH, RbH and CsH, but not Li2, as the system is too small (in the STO-6G basis) to carry out a $\mathcal{\kappa}$ versus $m$ analysis. We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase $n_{v}$. Figure 6 presents our results. We see that the condition number varies as a polylogarithmic function of system size. This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A14 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for KH, RbH, and CsH. The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix. We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock (HF) configuration. The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the HF configuration). Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value. Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size.
@@
Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches, via PFDs versus number of shots (denoted in the figure as ‘No. of shots’), for two representative molecules, Li2 (sub-figure (a)) and KH (sub-figure (b)). The centre of each unfilled circle represents the average PFD over a range of 30 repetitions. Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40). The zero PFD value is marked with a horizontal line for ease of visualization.
We also observe in each sub-figure (that is, for a fixed number of occupied orbitals) that for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH. This trend can be explained qualitatively using chemical intuition, by looking at the considered orbitals for the (6, 6) case as a representative example. Recalling that each eigenvalue of $A$ can be thought of as an excited state energy, the highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively. It is clear that $2p$ of K has much lower orbital energy than the $3d$ orbital of Rb. Thus, we can expect that the $2p$longrightarrow virtual excitation in KH has a higher excitation energy associated with it than the 3d $\longrightarrow$ virtual excitation in RbH. Similarly, we can expect that the $4d$longrightarrow virtual excited state of CsH is lower in energy than the $3d$longrightarrow virtual excited state of RbH. In summary, we can expect that the excitation energies decrease monotonically as we go from KH to CsH. We combine this observation with our data from Table A14 where the maximum eigenvalue grows much faster than the decrease in minimum eigenvalue from KH to RbH to CsH for a given choice of $n_{o}$ and $n_{v}$, to infer that $\mathcal{\kappa}$ decreases monotonically from KH to CsH.
Furthermore, in Figure 6(c), KH’s condition number reaches a staggering 4000 with all the virtuals considered, whereas in Figures 6(a) and (b), $\mathcal{\kappa}$ is substantially lower. This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals. Since KH has only 10 occupied orbitals, the active space contains the $1s$ orbital of K. We can invoke the fact that the orbital energy gap is larger between the inner s orbitals, for example, and the density of orbitals becomes higher in the vicinity of highest occupied molecular orbital and the lowest unoccupied molecular orbital. Thus, if we include inner core orbitals in our active space, we expect the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger. In contrast, for the 6, 8, and the 10 occupied orbitals’ cases, RbH and CsH have comparable condition numbers. This can be attributed to the fact that the lowest orbital is always the $3d$ of Rb in all the three cases. Since the orbital character of the lowest orbital is the same among the three active spaces, they yield very similar condition numbers.
@@
We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may be a reasonable direction to consider. However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation.
@@ III.2.3 HHL and Psi-HHL results
We begin by noting that it is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon. We thus move to Figure 7 (with the accompanying data given in Table A15 of the Appendix), where we carry out a study on the same lines as Figure 5. We vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions. For this analysis, we consider Li2 and KH. The details are:
@@
We pick these two systems, since in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300. Repeating the calculations for RbH and CsH may not give any new insights, and hence we restrict the results only to Li2 and KH. For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots. The data clearly shows that for Li2, the Psi-HHL approach always has a smaller spread in PFDs when compared to HHL. For KH, we observe that at 1000 shots, HHL fails, while Psi-HHL does not, but comes with large error bars. As we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller, but Psi-HHL consistently has smaller error bars than HHL.
@@
We now move to our next set of results, namely the PEC for the KH molecule. Since it is computationally not feasible to pick the larger system sizes from Figure 6, we choose the 6 occupied and 6 virtual orbitals case. This choice corresponds to a 26-qubit calculation with an $A$ matrix size of $256\times 256$. We consider geometries in the neighbourhood of the equilibrium bond length (3 geometries in steps of 0.47 Bohr on either side of the equilibrium bond length). With such a choice, the correlation energies required to be captured ranges from about 0.3 mHa (milli-Hartree) for 5.65 Bohr geometry to 3 mHa for the 2.82 Bohr geometry (the two extreme points in the 7 geometries considered), thus allowing us to check the performance of HHL and Psi-HHL across a range of correlation energy values. Calculating energies across a range of bond lengths further allows us to check across a (limited) range of condition numbers (the smallest being 365.40 and the largest 393.21) the behaviour of HHL and Psi-HHL. Each calculation involves 50000 shots and our reported results are averaged over 30 repetitions. For our Psi-HHL computations, we pick $\alpha$ of 60 degrees. Figure 8(a) (and the accompanying Table A16 of the Appendix) shows the results for $\Delta E$, which is the difference between HHL (Psi-HHL) total energy and and the classical LCC value for the considered geometries. We find that Psi-HHL performs better than HHL in predicting correlation energies, especially on the extreme ends of our PEC when $\mathcal{\kappa}$ values are relatively larger than for those points around the equilibrium geometry. When we examine Figure 8(b), we see that Psi-HHL performs consistently better than HHL. Furthermore, the spread in results across 30 repetitions follows a trend globally: it increases from 2.82 through 5.65 Bohr calculations. We recall that the correlation energies themselves decrease monotonically and by as much as an order of magnitude as we move from 2.82 Bohr to 5.65 Bohr geometries. Thus, although the condition numbers are comparable (for example, 374.93 and 371.38 for the 3.29 and the 4.71 Bohr cases respectively), since the quantity to be captured (the correlation energy) itself is much smaller for the latter, we need to supply more shots to restrict the spread in PFD across repeats. However, since we keep the number of shots fixed throughout the PEC, we observe a larger spread in PFDs. We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using the HOM circuit, with 1000 shots and average each of our results over 20 repetitions. $a$ is varied from 0.26 to 0.99 (below 0.26, we get ‘nan’ that is, situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result). We observe the same trend as in Figure 8(b) (see Figure A12 of Appendix).
Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs versus bond lengths (sub-figure (b)) of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots). The associated value of $\mathcal{\kappa}$ is shown above or below each data point.